# Iwahori Matsumoto presentation for modules of Iwahori fixed functions on symmetric spaces

### v2

@@ 7 Determining the structure constants $\gamma_{f,g}$
In this section we focus on the calculation of the constants $\gamma_{f,g}$ occurring in the formula
$(T_{s}+1)1_{x}=\gamma_{f,g}D_{f,o}$
proved in Proposition 4.3.
Let $s\in\tilde{\Delta}$ be a simple reflection. Let $\mathcal{C}$ be the chamber fixed by $I$ and let $g\in G$ be a representative of the orbit $x\in I\backslash G/H.$ Let $\mathcal{A}$ be a $\sigma$ stable apartment that contains $g\mathcal{C}$ and $gs\mathcal{C}$.
Consider the face $f=g\mathcal{C}\cap gs\mathcal{C}$.
In the case $l_{\sigma}(s\times x)\neq l_{\sigma}(x)$ we know the value of $\gamma_{f,g}$, its $1$ if $l_{\sigma}(s\times x)>l_{\sigma}(x)$ and $q$ if $l_{\sigma}(s\times x)<l_{\sigma}(x)$.
We assume that $l_{\sigma}(s\times x)=l_{\sigma}(x)$. This implies that the affine hyper plane spanned by $f$ is $\sigma$ stable (this was explained during the proof of Lemma 5.5). Thus $f$ satisfies the assumption of the previous section.
Let $f^{\prime}=f\cup\sigma(f)$. Denote by $C_{f}$ the set of all chambers in $\mathcal{B}_{G}$ that contain $f$. Let $G_{f^{\prime}}$ be the reductive quotient associated to $f^{\prime}$, it is a reductive group over the residue field $k$. $\sigma$ acts on $G_{f^{\prime}}$ as was shown in the previous section. Let $B(G_{f^{\prime}})$ be the flag variety of $G_{f^{\prime}}$.
@@
By the definition $\gamma_{f,g}$ is equal to the number of chambers in $C_{f}$ in the $H\cap Ker(\omega)$ orbit of $g\mathcal{C}$. By Proposition 6.6 this is equal to the size of the $G^{\sigma}_{f}$ orbit of $b_{f}(g\mathcal{C})$ in the flag variety of $G_{f}$. ∎
@@
By proposition 2.25 the root system of $G_{f^{\prime}}$ consists of the roots constant on $f$, $f$ is a face of co-dimension 1 so there is a single root (up to a sign) in the root system of $G_{f^{\prime}}$. Therefore the semisimple rank of $G_{f^{\prime}}$ is 1. ∎
Thus, in order to compute the value of $\gamma_{f,g}$ we need to compute the size of an orbit of a symmetric subgroup of a reductive group of semisimple rank 1, on the flag variety of this reductive group.
@@
Let $K$ be the $k$ points of one of the groups from the previous proposition, let $\sigma$ be an algebraic involution on $K$ and let $M=K^{\sigma}$. Let $Z=Z(K)^{0}$ be the connected center of $K$.
$\sigma$ induces an involution on $K/Z$, and $M$ projects to fixed points of this involution on $K/Z$.
Notice that the action of $K$ on its flag variety factors through $K/Z$. In the first case of proposition 7.3 we have $K/Z=SL_{2}(k)$, in the two other cases we have $K/Z=PGL_{2}(k)$. Thus, we are interested in symmetric subgroups of $PGL_{2}(k),SL_{2}(k)$.
@@
    $\tilde{T}$ a non split torus.
@@
In the third case this might not be true. If we restrict to $SL_{2}\subset GL_{2}\times\mathbb{G}^{r}_{m}$ it will be true, meaning all $\sigma$ stable elements of $SL_{2}/Z$ do lift to $SL_{2}(k)\cap M$. This means that $M/Z$ must contain the index 4 subgroup $T^{2}\subset N_{PGL_{2}(k)}(T)$ both in the case of a split and in the case of non split torus.
Consider $\tilde{M}=Im(M\rightarrow Aut(\mathbb{P}^{1}))$ and its orbits on the flag variety of $K$, the options are summarized by the next proposition:
@@
Let $\tilde{M}$ be the image of $M$ in the automoprhism group of the flag variety of $K$, which is isomorphic to $PGL_{2}(k)$.
Let $T_{s}$ be a split torus of $PGL_{2}(K)$ and let $w\in N_{PGL_{2}(k)}(T_{s})$ be an element that normalizes $T_{s}$ but isn’t in $T_{s}$.
Let $T_{n}$ be a non split torus of $PGL_{2}(K)$ and let $\tilde{w}\in N_{PGL_{2}(k)}(T_{n})$ be an element that normalizes $T_{n}$ but isn’t in $T_{n}$.
The options for $\tilde{M}$ up to conjugation, as well as the options for the numbers and sizes of its orbits on $\mathbb{P}^{1}$ are given by the following table.
@@
$T_{s}$ 3 $11,1,q-$ 1
$T_{s}^{2}$ 4 $1,1,\frac{q-1}{2},\frac{q-1}{2}$
$T_{s}\cup wT_{s}$ 2 $12,q-$ 1
$wT_{s}^{2}\cup T_{s}^{2}$ 3 $2,\frac{q-1}{2},\frac{q-1}{2}$
$wT_{s}\cup T_{s}^{2}$ 2 $12,q-$ 1
@@
This is follows from a simple computation in each of the case in the previous proposition, and in the case $K=GL_{2}(k)\times(k^{\times})^{r}$ and $T^{2}\subset\tilde{M}\subset N_{PGL_{2}(k)}(T)$ where $T$ might be a split or a non split torus. ∎
@@
The possible values for $\gamma_{f,g}$ are $11,2,\frac{q-1}{2},q-1,q,q+$ 1 and $\frac{q+1}{2}$.
@@
By theorem 5.3 if $l_{\sigma}(gs\mathcal{C})\neq l_{\sigma}(g\mathcal{C})$ then $\gamma_{f,g}$ is either $1$ or $q$. By proposition 7.5 if $l_{\sigma}(gs\mathcal{C})=l_{\sigma}(g\mathcal{C})$ then $\gamma_{f,g}$ is one of $11,2,\frac{q-1}{2},q-1,q+$ 1 and $\frac{q+1}{2}$. ∎
@@
All the values mentioned in corollary 7.6 can be obtained.
The values of $11,q,q+$ 1 can be obtained in the case of $G=SL_{4}$ and $H=Sp_{4}$.
The value $\frac{q-1}{2}$ can be obtained in the case $G=SL_{2}$, $H=T$ a maximal split torus.
The value $q-1$ can be obtained in the case $G=GL_{2}$ and $H=T$ a maximal split torus.
The value $2$ can be obtained in the case $G=PGL_{2}$ and $H=N_{G}(T)$ normalizer of a maximal split torus.
The value $\frac{q+1}{2}$ can be obtained in the case $G=SL_{2}$ and $H=\tilde{T}$ a non split torus.
@@
Let $\mathcal{C}_{1},...\mathcal{C}_{q+1}$ be the chambers that contain $f$. The set
$\{l_{\sigma}(\mathcal{C}_{1}),...,l_{\sigma}(\mathcal{C}_{q+1})\}$ is of size at most 2. Furthermore, the difference between different $l_{\sigma}$ lengths is at most 2 if $l_{\sigma}(x)\neq l_{\sigma}(s\times x)$ and at most 1 if $l_{\sigma}(x)=l_{\sigma}(s\times x)$.
@@
If $l_{\sigma}(x)<l_{\sigma}(s\times x)$ then by lemma 5.5 every chamber in $\{\mathcal{C}_{1},...\mathcal{C}_{q+1}\}$ which isn’t $g\mathcal{C}$ is contained in a $\sigma$ stable apartment, $\mathcal{A}$, that satisfies $g\mathcal{C}\subset\mathcal{A}$. Thus by Lemma 3.1 all chambers in $\{\mathcal{C}_{1},...\mathcal{C}_{q+1}\}$ which are different from $g\mathcal{C}$ are in the same $H$ orbit and hence have the same $l_{\sigma}$ length. This $l_{\sigma}$ length is equal to $l_{\sigma}(s\times x)=l_{\sigma}(x)+2$. the same argument also works in the case $l_{\sigma}(x)>l_{\sigma}(s\times x)$.
Now assume that $l_{\sigma}(x)=l_{\sigma}(s\times x)$. Let $\mathcal{C}^{\prime}$ be a chamber that contains $f$ and let $\mathcal{C}^{\prime\prime}$ be a chamber that contains $\sigma(f)$. There is an apartment $\mathcal{A}$ that contains both $\mathcal{C}^{\prime}$ and $\mathcal{C}^{\prime\prime}$. Denote by $\overline{f}$ the affine subspace spanned by $f$ inside $\mathcal{A}$. There are two cases, either $\mathcal{C}^{\prime}$ and $\mathcal{C}^{\prime\prime}$ are on the same side of $\overline{f}$ or they are on different sides. This distinction determines $d(\mathcal{C}^{\prime},\mathcal{C}^{\prime\prime})$, in the case they are on different sides their distance is larger by $1$ than the distance in the case they are on the same side.
For any chamber $\mathcal{C}^{\prime}$ that contains $f$, $\sigma(\mathcal{C}^{\prime})$ contains $\sigma(f)$. Therefore there are two possible values for $l_{\sigma}(\mathcal{C}^{\prime})=d(\mathcal{C}^{\prime},\sigma(\mathcal{C}^{\prime}))$ and the difference between them is $1$
@@
Notice that we proved that if $l_{\sigma}(x)=l_{\sigma}(s\times x)$ then $l_{\sigma}(g\mathcal{C})$ might have one of two values, those values depend only on $f$. $l_{\sigma}(g\mathcal{C})$ will obtain the minimal value if and only if $\sigma(g\mathcal{C})$ and $g\mathcal{C}$ are on the same side of $f$ inside some apartment.
@@
Assume $f$ is a face of codimension 1, let $f^{\prime}=f\cup\sigma(f)$ and let $G_{f^{\prime}}$ be the reductive group of semisimple rank 1 defined over $k$ associated to $f^{\prime}$. Recall we have an action of $\sigma$ on $G_{f^{\prime}}$. Let $b_{f}$ be the map from $C_{f}$, the set of chambers that contain $f$, to $B(G_{f^{\prime}})$, the flag variety of $G_{f^{\prime}}$. Then $g\mathcal{C}$ and $\sigma(g\mathcal{C})$ are on the same side of $f$ in some apartment if and only if $b_{L}(g\mathcal{C})$ is $\sigma$ stable.
@@
[…]
Let $p\in B(G_{f^{\prime}})$, $p$ is fixed by $\sigma$ if and only if $\sigma(res^{f^{\prime}}_{f}(p))=res^{f^{\prime}}_{\sigma(f)}(p)$. Let $\mathcal{C}_{p}\in C_{f}$ be the chamber that corresponds to $res^{f^{\prime}}_{f}(p)$ and let $\mathcal{C}_{p}^{\prime}\in C_{\sigma(f)}$ the chamber that corresponds to $res^{f^{\prime}}_{\sigma(f)}(p)$. We know that $\sigma(p)=p$ if and only if $\sigma(\mathcal{C}_{p})=\mathcal{C}_{p}^{\prime}$, thus it is enough to prove that $\mathcal{C}_{p}$ and $\mathcal{C}_{p}^{\prime}$ are always on the same side of $f$.
We can lift $p$ to a subgroup of $P_{f^{\prime}}$, we denote it by $S$. $S\subset P_{f^{\prime}}={\cal G}_{f^{\prime}}(\mathcal{O})$. The restriction maps are just the restrictions to the groups $P_{f}$ and $P_{\sigma(f)}$ that contain $P_{f^{\prime}}$ and the passage to chambers in $C_{f}$ ($C_{\sigma(f)}$) is by taking the points in the building fixed by $S$ and finding inside them a chamber that contains $f$ ($\sigma(f)$). Therefore we only need to show that if $\mathcal{C}^{\prime}$ is a chamber that contains $f$ and $\mathcal{A}$ is an apartment that contains both $\mathcal{C}^{\prime}$ and $\sigma(f)$, and $g^{\prime}\in S$ fixes both $\mathcal{C}^{\prime}$ and $\sigma(f)$ pointwise then $g^{\prime}$ fixes a chamber in $\mathcal{A}$ that contains $\sigma(f)$ on the same side of $f$ as $\mathcal{C}$. This follows from the fact that the set of elements in $\mathcal{A}$ fixed by $g^{\prime}$ is convex (see Theorem 2.16 and part 5 of Theorem 2.13). ∎
@@
Assume $l_{\sigma}(x)=l_{\sigma}(s\times x)$ and that not all of $l_{\sigma}(\mathcal{C}_{1}),...,l_{\sigma}(\mathcal{C}_{q+1})$ are equal. Let $b_{f}$ be the map from $C_{f}$, the set of chambers that contain $f$, to $B(G_{f^{\prime}})$ the flag variety of $G_{f^{\prime}}$. Let $\mathcal{C}_{i}$ be a chamber that contains $f$. $l_{\sigma}(\mathcal{C}_{i})$ is minimal among $l_{\sigma}(\mathcal{C}_{1}),...,l_{\sigma}(\mathcal{C}_{q+1})$ if and only if $b_{f}(\mathcal{C}_{i})$ is $\sigma$ stable.
@@
This follows immediately from the previous proposition and the remark before it. ∎
Using the previous result we can determine $\gamma_{f,g}$ using the $l_{\sigma}$ lengths of the orbits of the chambers that contain $f$.
As explained above, the calculation of $T_{s}1_{x}$ was reduced to the calculation of $\gamma_{f,g}$. To determine this number we need to know further information regarding the lengths of certain chambers. Let $\mathcal{C}_{1},...,\mathcal{C}_{q+1}$ be all the chambers that contain $f$, Let $\mathcal{O}_{f}$ be the set of their $H$ orbits. By proposition 7.8 there are at most two possibilities for the $l_{\sigma}$ lengths of orbits in $\mathcal{O}_{f}$. Denote by $n_{max},n_{min}$ be the number of orbits in $\mathcal{O}_{f}$ of maximal/minimal $l_{\sigma}$ length respectively. Let $\delta_{x,max}$ be 1 if $l_{\sigma}(g\mathcal{C})$ is maximal among $l_{\sigma}(\mathcal{C}_{1}),...,l_{\sigma}(\mathcal{C}_{q+1})$ and 00 otherwise. $\delta_{x,min}$ is defined similarly.
@@
We have $n_{max},n_{min}\in\{1,2\}$. If $l_{\sigma}(x)=l_{\sigma}(s\times x)$ then
@@
From Propositions 7.11, 6.3 and 6.4 we know that the number of orbits of minimal/maximal $\sigma$ length is equal to the number of $G^{\sigma}_{f}$ orbits on the flag variety of $G_{f^{\prime}}$ of minimal/maximal size. Now a simple check of the cases in Proposition 7.5 verifies the result.

### v3

@@ 7 Determining the structure constants $\gamma_{f,g}$
In this section, we focus on calculating the constants $\gamma_{f,g}$ occurring in the formula
$1_{x}(T_{s}+1)=\gamma_{f,g}D_{f,o}$
as proved in Proposition 4.3.
Let $s\in\tilde{\Delta}$ be a simple reflection. Let $\mathcal{C}$ be the chamber fixed by $I$ and let $g\in G$ be a representative of the orbit $x\in H\backslash G/I.$ Let $\mathcal{A}$ be a $\sigma$ stable apartment that contains $g\mathcal{C}$ and $gs\mathcal{C}$.
Consider the facet $f=g\mathcal{C}\cap gs\mathcal{C}$.
In the case $l_{\sigma}(x\times s)\neq l_{\sigma}(x)$ we know the value of $\gamma_{f,g}$. It equals $1$ if $l_{\sigma}(x\times s)>l_{\sigma}(x)$ and equals $q$ if $l_{\sigma}(x\times s)<l_{\sigma}(x)$.
We now assume that $l_{\sigma}(x\times s)=l_{\sigma}(x)$. This implies that the affine hyper plane spanned by $f$ is $\sigma$ stable (this was explained during the proof of Lemma 5.5). Thus, $f$ satisfies the assumption of the previous section.
Let $f^{\prime}=f\cup\sigma(f)$. Denote by $C_{f}$ the set of all chambers in $\mathcal{B}_{G}$ that contain $f$. Let $G_{f^{\prime}}$ be the reductive quotient associated to $f^{\prime}$, it is a reductive group over the residue field $k$. The map $\sigma$ acts on $G_{f^{\prime}}$ as was shown in the previous section. Let $B(G_{f^{\prime}})$ be the flag variety of $G_{f^{\prime}}$.
@@
By definition, $\gamma_{f,g}$ is equal to the number of chambers in $C_{f}$ in the $H\cap Ker(\omega)$ orbit of $g\mathcal{C}$. By Proposition 6.6 this is equal to the size of the $G^{\sigma}_{f}$ orbit of $b_{f}(g\mathcal{C})$ in the flag variety of $G_{f}$. ∎
@@
By Proposition 2.26 the root system of $G_{f^{\prime}}$ consists of the roots constant on $f$, $f$ is a facet of codimension 1 so there is a single root (up to a sign) in the root system of $G_{f^{\prime}}$. Therefore the semisimple rank of $G_{f^{\prime}}$ is 1. ∎
Example 7.3.
In Example 6.7 we saw several examples of this when $L$ had codimension 1. In the first example we had $G_{f^{\prime}}\cong SL_{3}\cap GL_{2}\times GL_{1}$ and in the second example we had $G_{f^{\prime}}\cong SL_{2}\times GL_{1}$. In both examples $G_{f^{\prime}}$ has semisimple rank 1.
Thus, to compute the value of $\gamma_{f,g}$ we need to compute the size of an orbit of a symmetric subgroup of a split reductive group of semisimple rank 1, on the flag variety of this reductive group.
@@
Let $K$ be the $k$ points of one of the groups from the previous proposition, let $\sigma$ be an algebraic involution on $K$ and let $M=K^{\sigma}$. Let $Z=Z(K)^{0}$ be the connected component of the center of $K$.
$\sigma$ induces an involution on $K/Z$, and $M$ projects into the fixed points of this involution on $K/Z$.
Note that the action of $K$ on its flag variety factors through $K/Z$. In the first case of Proposition 7.4 we have $K/Z=SL_{2}(k)$, in the two other cases we have $K/Z=PGL_{2}(k)$. Thus, we are interested in symmetric subgroups of $PGL_{2}(k),SL_{2}(k)$.
@@
    $\tilde{T}$ a non-split torus.
@@
In the third case this might not be true. However, if we restrict to $SL_{2}\subset GL_{2}\times\mathbb{G}^{r}_{m}$ it is true. That is, all $\sigma$ stable elements of $SL_{2}/Z$ do lift to $SL_{2}\cap M$. This means that $M/Z$ must contain the index 4 subgroup $T^{2}\subset N_{PGL_{2}(k)}(T)$ both in the case of a split and in the case of non-split torus.
Consider $\tilde{M}=Im(M\rightarrow Aut(\mathbb{P}^{1}))$ and its orbits on the flag variety of $K$, the possibilities are summarized by the next proposition:
@@
Let $\tilde{M}$ be the image of $M$ in the automorphism group of the flag variety of $K$, which is isomorphic to $PGL_{2}(k)$.
Let $T_{sp}$ be a split torus of $PGL_{2}(k)$ and let $w\in N_{PGL_{2}(k)}(T_{sp})$ be an element that normalizes $T_{sp}$ but is not in $T_{sp}$.
Let $T_{n}$ be a non-split torus of $PGL_{2}(k)$ and let $\tilde{w}\in N_{PGL_{2}(k)}(T_{n})$ be an element that normalizes $T_{n}$ but is not in $T_{n}$.
The options for $\tilde{M}$ up to conjugation, as well as the options for the numbers and sizes of its orbits on $\mathbb{P}^{1}$ are given by Table 1.
Table 1: Subgroups $\tilde{M}$ of $PGL_{2}(k)$ which can be obtained as the image of $M$. Together with the number and sizes of their orbits on $\mathbb{P}^{1}$.
@@
$T_{sp}$ 3 1,1,q11,1,q-1
$T_{sp}^{2}$ 4 $1,1,\frac{q-1}{2},\frac{q-1}{2}$
$T_{sp}\cup wT_{sp}$ 2 2,q12,q-1
$wT_{sp}^{2}\cup T_{sp}^{2}$ 3 $2,\frac{q-1}{2},\frac{q-1}{2}$
$wT_{sp}\cup T_{sp}^{2}$ 2 2,q12,q-1
@@
This follows from a simple computation in each case given by the previous proposition. As well as in the case $K=GL_{2}(k)\times(k^{\times})^{r}$ and $T^{2}\subset\tilde{M}\subset N_{PGL_{2}(k)}(T)$ for $T$ a torus which may be either split or a non-split. ∎
@@
The possible values of $\gamma_{f,g}$ are $1,2,\frac{q-1}{2},q-1,q,q+1$ and $\frac{q+1}{2}$.
@@
By Theorem 5.3 if $l_{\sigma}(gs\mathcal{C})\neq l_{\sigma}(g\mathcal{C})$ then $\gamma_{f,g}$ is either 11 or $q$. By Proposition 7.6 if $l_{\sigma}(gs\mathcal{C})=l_{\sigma}(g\mathcal{C})$ then $\gamma_{f,g}$ is one of $1,2,\frac{q-1}{2},q-1,q+1$ and $\frac{q+1}{2}$. ∎
@@
All the values mentioned in Corollary 7.7 occur.
The values $1,q,q+1$ are obtained in the case $G=SL_{4}$ and $H=Sp_{4}$.
The value $\frac{q-1}{2}$ is obtained in the case $G=SL_{2}$, $H=T$ a maximal split torus.
The value q1q-1 is obtained in the case $G=GL_{2}$ and $H=T$ a maximal split torus.
The value $2$ is obtained in the case $G=PGL_{2}$ and $H=N_{G}(T)$ a normalizer of a maximal split torus.
The value $\frac{q+1}{2}$ is obtained in the case $G=SL_{2}$ and $H=\tilde{T}$ a non-split torus.
@@
Let $\mathcal{C}_{1},...\mathcal{C}_{q+1}$ be the collection of chambers that contain $f$. The set
$\{l_{\sigma}(\mathcal{C}_{1}),...,l_{\sigma}(\mathcal{C}_{q+1})\}$ is of size at most 2. Furthermore, the difference between different $l_{\sigma}$ lengths is at most 2 if $l_{\sigma}(x)\neq l_{\sigma}(x\times s)$ and at most 1 if $l_{\sigma}(x)=l_{\sigma}(x\times s)$.
@@
If $l_{\sigma}(x)<l_{\sigma}(x\times s)$ then by Lemma 5.5 every chamber in $\{\mathcal{C}_{1},...\mathcal{C}_{q+1}\}$ which is not equal to $g\mathcal{C}$ is contained in a $\sigma$ stable apartment, $\mathcal{A}$, that satisfies $g\mathcal{C}\subset\mathcal{A}$. Thus, by Proposition 3.1 all chambers in $\{\mathcal{C}_{1},...\mathcal{C}_{q+1}\}$ different from $g\mathcal{C}$ are in the same $H$ orbit and hence have the same $l_{\sigma}$ length. This $l_{\sigma}$ length is equal to $l_{\sigma}(x\times s)=l_{\sigma}(x)+2$. The same argument also works in the case $l_{\sigma}(x)>l_{\sigma}(x\times s)$.
Now assume that $l_{\sigma}(x)=l_{\sigma}(x\times s)$. Let $\mathcal{C}^{\prime}$ be a chamber that contains $f$ and let $\mathcal{C}^{\prime\prime}$ be a chamber that contains $\sigma(f)$. There is an apartment $\mathcal{A}$ that contains both $\mathcal{C}^{\prime}$ and $\mathcal{C}^{\prime\prime}$. Denote by $\overline{f}$ the affine subspace spanned by $f$ inside $\mathcal{A}$. There are two cases, either $\mathcal{C}^{\prime}$ and $\mathcal{C}^{\prime\prime}$ are on the same side of $\overline{f}$ or they are on different sides. This distinction determines $d(\mathcal{C}^{\prime},\mathcal{C}^{\prime\prime})$, in the case they are on different sides their distance is larger by $1$ than the distance in the case they are on the same side.
For any chamber $\mathcal{C}^{\prime}$ that contains $f$, $\sigma(\mathcal{C}^{\prime})$ contains $\sigma(f)$. Therefore there are two possible values for $l_{\sigma}(\mathcal{C}^{\prime})=d(\mathcal{C}^{\prime},\sigma(\mathcal{C}^{\prime}))$ and the difference between them is $1.$
@@
Notice that we proved that if $l_{\sigma}(x)=l_{\sigma}(x\times s)$ then $l_{\sigma}(g\mathcal{C})$ may take one of two values, those values depend only on $f$. $l_{\sigma}(g\mathcal{C})$ attains its minimal value if and only if $\sigma(g\mathcal{C})$ and $g\mathcal{C}$ are on the same side of $f$ inside some apartment.
@@
Assume $f$ is a facet of codimension 1, let $f^{\prime}=f\cup\sigma(f)$ and let $G_{f^{\prime}}$ be the reductive group of semisimple rank 1 defined over $k$ associated to $f^{\prime}$. Recall we have an action of $\sigma$ on $G_{f^{\prime}}$. Let $b_{f}$ be the map from $C_{f}$, the set of chambers that contain $f$, to $B(G_{f^{\prime}})$, the flag variety of $G_{f^{\prime}}$. Then $g\mathcal{C}$ and $\sigma(g\mathcal{C})$ are on the same side of $f$ in some apartment if and only if $b_{L}(g\mathcal{C})$ is $\sigma$ stable.
@@
$B(G_{f^{\prime}})$$B(G_{f^{\prime}})$$B(G_{f})$$B(G_{\sigma(f)})$$C_{f}$$C_{\sigma(f)}$$\scriptstyle{\sigma}$$\scriptstyle{res^{f^{\prime}}_{f}}$$\scriptstyle{res^{f^{\prime}}_{\sigma(f)}}$$\scriptstyle{\sigma}$$\scriptstyle{\overline{res}^{-1}_{f}}$$\scriptstyle{\overline{res}^{-1}_{\sigma(f)}}$$\scriptstyle{\sigma}$
Let $\in B(G_{f^{\prime}}), pp$ is fixed by $\sigma$ if and only if $\sigma(res^{f^{\prime}}_{f}(p))=res^{f^{\prime}}_{\sigma(f)}(p)$. Let $\mathcal{C}_{p}\in C_{f}$ be the chamber that corresponds to $res^{f^{\prime}}_{f}(p)$ and let $\mathcal{C}_{p}^{\prime}\in C_{\sigma(f)}$ the chamber that corresponds to $res^{f^{\prime}}_{\sigma(f)}(p)$. We know that $\sigma(p)=p$ if and only if $\sigma(\mathcal{C}_{p})=\mathcal{C}_{p}^{\prime}$, thus it is enough to prove that $\mathcal{C}_{p}$ and $\mathcal{C}_{p}^{\prime}$ are always on the same side of $f$.
We lift $p$ to a subgroup of $P_{f^{\prime}}$, we denote it by $S$. We have $S\subset P_{f^{\prime}}={\mathcal{G}}_{f^{\prime}}(\mathcal{O})$. The restriction maps are just the restriction to the groups $P_{f}$ and $P_{\sigma(f)}$ that contain $P_{f^{\prime}}$. The passage to chambers in $C_{f}$ (or $C_{\sigma(f)}$) is given by taking the points in the building fixed by $S$ and finding inside them a chamber that contains $f$ (respectively $\sigma(f)$). Therefore, we only need to show that if $\mathcal{C}^{\prime}$ is a chamber that contains $f$, $\mathcal{A}$ is an apartment that contains both $\mathcal{C}^{\prime}$ and $\sigma(f)$, and $g^{\prime}\in S$ fixes both $\mathcal{C}^{\prime}$ and $\sigma(f)$ pointwise then $g^{\prime}$ fixes a chamber in $\mathcal{A}$ that contains $\sigma(f)$ on the same side of $f$ as $\mathcal{C}$. This follows from the fact that the set of elements in $\mathcal{A}$ fixed by $g^{\prime}$ is convex (see Theorem 2.17 and part 5 of Theorem 2.13). ∎
@@
Assume $l_{\sigma}(x)=l_{\sigma}(x\times s)$ and that not all of $l_{\sigma}(\mathcal{C}_{1}),...,l_{\sigma}(\mathcal{C}_{q+1})$ are equal. Let $b_{f}$ be the map from $C_{f}$, the set of chambers that contain $f$, to $B(G_{f^{\prime}})$ the flag variety of $G_{f^{\prime}}$. Let $\mathcal{C}_{i}$ be a chamber that contains $f$. The value $l_{\sigma}(\mathcal{C}_{i})$ is minimal among $l_{\sigma}(\mathcal{C}_{1}),...,l_{\sigma}(\mathcal{C}_{q+1})$ if and only if $b_{f}(\mathcal{C}_{i})$ is $\sigma$ stable.
@@
This follows immediately from the previous proposition and Remark 7.10. ∎
Using the previous result we can determine $\gamma_{f,g}$ from the $l_{\sigma}$ lengths of the orbits of the chambers that contain $f$.
As explained above, the calculation of $T_{s}1_{x}$ was reduced to the calculation of $\gamma_{f,g}$. To determine this number we need to know further information regarding the lengths of certain chambers. Let $\mathcal{C}_{1},...,\mathcal{C}_{q+1}$ be all the chambers that contain $f$. Let $\mathcal{O}_{f}$ be the set of their $H$ orbits. By Proposition 7.9 there are at most two possibilities for the $l_{\sigma}$ lengths of orbits in $\mathcal{O}_{f}$. Denote by $n_{max},n_{min}$ be the number of orbits in $\mathcal{O}_{f}$ of maximal/minimal $l_{\sigma}$ length respectively. Let $\delta_{x,max}$ be 1 if $l_{\sigma}(g\mathcal{C})$ is maximal among $l_{\sigma}(\mathcal{C}_{1}),...,l_{\sigma}(\mathcal{C}_{q+1})$ and 0 otherwise. The value $\delta_{x,min}$ is defined in a similar way.
@@
We have $n_{max},n_{min}\in\{1,2\}$. If $l_{\sigma}(x)=l_{\sigma}(x\times s)$ then
@@
By Propositions 7.12, 6.3 and 6.4 we know that the number of orbits of minimal/maximal $\sigma$ length is equal to the number of $G^{\sigma}_{f}$ orbits on the flag variety of $G_{f^{\prime}}$ of minimal/maximal size. Now a simple check of the cases in Proposition 7.6 verifies the result.